# Empowering Snapshot Compressive Imaging: Spatial-Spectral State Space Model with Across-Scanning and Local Enhancement

### v1

@@ 3.1 Architecture of ASLE-SSM
The overall architecture of our ASLE-SSM is depicted in Fig. 2. ASLE-SSM adopts a U-shaped network with the spatial-spectral SSM block as the basic unit. The structure consists of the encoder, bottleneck, and decoder. Given the vectorized measurement $\mathbf{y}\in\mathbb{R}^{H(W+d(N_{\lambda}-1))}$, we reverse the dispersion and shift back $\mathbf{Y}$ to initialize the input signal $\mathbf{X}\in\mathbb{R}^{H\times W\times N_{\lambda}}$:
@@
We first concatenate the input $\mathbf{X_{0}}$ with the physical mask $\mathbf{Y}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ and then employ a 1$\times$1 convolution layer to integrate information from the original signal and the physical mask. Moreover, the integrated input $\mathbf{X_{0}}$ undergoes a 3$\times$3 convolution embedding into a feature $\mathbf{X_{0}^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times C}$. Subsequently, $\mathbf{X_{0}^{{}^{\prime}}}$ passes through the downsample and upsample process. After passing through the decoder, we employ a 3$\times$3 convolution mapping to obtain the residual information. By computing the sum of the integrated input and residual output, we finally obtain the reconstructed HSIs. In our implementation, we change the number of spatial-spectral SSM blocks in the encoder, bottleneck, and decoder to realize the small, middle, and large sizes of ASLE-SSM models.
@@
     Algorithms   Params   GFLOPs      S1      S2      S3      S4      S5      S6      S7      S8      S9      S10     Avg
GAP-TV gap_tv
26.04
0.817
21.66
0.724
24.86
0.732
30.51
0.875
24.33
0.778
25.11
0.790
18.28
0.730
23.94
0.780
21.77
0.732
23.08
0.721
23.96
0.768
DeSCI desci
28.38
0.803
26.00
0.701
23.11
0.730
28.26
0.855
25.41
0.778
24.66
0.764
24.96
0.725
24.15
0.747
23.56
0.701
24.17
0.677
25.27
0.748
$\lambda$-Net lambda
62.64M
117.98
30.10
0.849
28.49
0.805
27.73
0.870
37.01
0.934
26.19
0.817
28.64
0.853
26.47
0.806
26.09
0.831
27.50
0.826
27.13
0.816
28.53
0.841
TSA-Net tsa_net
44.25M
110.06
32.31
0.894
31.03
0.863
32.15
0.916
37.95
0.958
29.47
0.884
31.06
0.902
30.02
0.880
29.22
0.886
31.14
0.909
29.18
0.861
31.35
0.895
HDNet hdnet
2.37M
154.76
34.96
0.937
35.64
0.943
35.55
0.946
41.64
0.976
32.56
0.948
34.33
0.954
33.27
0.928
32.26
0.945
34.17
0.944
32.22
0.940
34.66
0.946
DIP-HSI self
@@
DiffSCI diffsci
@@
ASLE-SSM-S
0.81M
7.01
35.01
0.938
34.57
0.932
36.19
0.953
42.04
@@
32.40
0.943
33.99
0.949
33.42
0.928
32.38
0.945
@@
33.03
0.956
34.92
0.963
34.26
0.940
33.47
0.961
36.08
0.962
32.56
0.950
35.74
0.958
3.2 Local Enhancement SSM (LE-SSM) Mechanism
Although Mamba is effective in modeling long-range dependencies, directly using the SSM mechanism often leads to less effective in extracting local features. Recent models tend to incorporate CNNs into the network structure, utilizing their local feature extraction capabilities to improve the performance of original SSMs. However, such methods would increase the parameter count, diminishing the lightweight advantage of the original SSM mechanism. While directly adding convolutional layers can enhance the model’s performance, it does not fundamentally solve the problem.
Inspired by LocalMamba localmamba , we introduce a simplified local mechanism in the spectral dimension. We use two scanning directions as the global branch, scanning the image row by row in both forward and backward directions while using the other two scanning directions as local branches to unfold and scan local patches. The specific details of the scanning process are illustrated in Fig. 2(d). Given the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, we first generate four copies of input data with different arrangements. For the global branch, we directly flatten the feature map into $\mathbf{X_{1}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, and then reverse it to generate $\mathbf{X_{2}}$. For the local branch, we first set the size of our patch flattening each patch into a 1D sequence, and then linking each patch together to generate $\mathbf{X_{3}}$. Similarly, we reverse $\mathbf{X_{3}}$ to obtain $\mathbf{X_{4}}$. After the generation of input data in four directions, we conduct the process as follows:
$\mathbf{\hat{X}}=LE\text{-}Spatial\text{-}SSM(\mathbf{X^{{}^{\prime}}}).$
The structure of LE-Spatial-SSM is shown in Fig. 2(b). It is worth noting that our method does not employ any CNNs or spatial and channel attention mechanisms based on MLPs, ensuring that our method does not add any extra parameters or time consumption compared to the global cross-scan mechanism liu2024vmamba . However, it also means we need to carefully choose the size of local patches, or else the reconstruction result will even become worse. In our implementation, we use patches of size 4$\times$4 to achieve a relatively better performance.
3.3 Across-Scanning SSM (AS-SSM) Mechanism
Our LE-SSM mechanism establishes a local-enhanced global receptive field but does not fully exploit the rich spectral information inherent in HSIs, leading to suboptimal reconstruction. Recently, some methods cumamba ; s2mamba have utilized and integrated channel information by directly scanning along the channel dimension (spectral dimension), but this scanning method does not consider the local features in both channel and spatial dimensions. In other words, neighboring pixels and bands often have certain similarities, but scanning the spectral dimension directly may lead to a distance increase between data points that are highly correlated.
To address the above issues, we propose a local scanning method that integrates local features from both spatial and spectral dimensions. This mechanism attempts to exploit the local properties of HSIs in both spatial and spectral dimensions while leveraging the rich spectral information to guide the reconstruction process. First, we rearrange the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ into patches along the spatial context by dividing the input HSIs into a series of smaller data cubes $\mathbf{S}\in\mathbb{R}^{P\times P\times C}$, which will be flattened into sequences for scanning. After that, we organize the scanning order within each sequence based on our spatial-spectral local cubes $\mathbf{C}\in\mathbb{R}^{h\times w\times c}$. As shown in Fig. 2(e), different colors indicate each spatial-spectral local cube. The scanning order of each data point is also depicted. The process can be formulated as:
$\mathbf{C}=Divide(\mathbf{X}),$ (8)
$\mathbf{C^{{}^{\prime}}}=AS\text{-}Spectral\text{-}Spatial\text{-}SSM(\mathbf{C}),$
where $\mathbf{C^{{}^{\prime}}}$ denotes the output patch of AS-SSM. After the SSM mechanism, we reorganize the patches back into the same order as the input $\mathbf{X}$. By scanning the data points within each data cube, we can leverage the local similarity between adjacent spectral bands and pixels, thus achieving better results than previous methods that directly scan along the spectral dimension.
@@ 4.1 Experimental Settings
The effectiveness of our ASLE-SSM network is verified on both simulation and real datasets, by using 28 wavelengths ranging from 450nm to 650nm, acquired via spectral interpolation manipulation.
Simulation HSI Data. Two simulated HSI datasets CAVE cave and KAIST kaist are utilized to conduct the simulation experiment. The CAVE dataset contains 32 HSIs with a spatial size of $512\times 512$, while the KAIST dataset includes 30 HSIs with a spatial size of $2704\times 3376$. Following the experimental settings of TSA-Net tsa_net , we designate CAVE as the training set and select 10 scenes from the KAIST as the testing set.
Real HSI Data. We adopt the real HSIs obtained by the CASSI system, as developed in TSA-Net tsa_net for testing.
Evaluation Metrics. We select peak signal-to-noise ratio (PSNR) PSNR and structural similarity (SSIM) ssim as the evaluation metrics to assess the quality of HSI reconstruction.
Implementation Details. Our ASLE-SSM is implemented using PyTorch and we conduct model training with Adam optimizer ($\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999) for 300 epochs. The learning rate is set to $4\times 10^{-4}$ at the beginning and is halved every 50 epochs during the training procedure. Image patches with size $256\times 256$ are randomly cropped from the 3D HSI cubes with 28 channels for training in the simulation and real experiments. The shifting step $d$ in the dispersion process is set to 2 and the batch size is 2. We apply the data augmentation techniques including random flipping and rotation. The models are trained on an RTX 4090 GPU.
Refer to caption Scene 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Meas GT Patch ADMM-Net CST-L+ DGSMP DIP-HSI HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DiffSCI $\lambda$-Net MST-L PnP-CASSI TSA-Net ASLE-SSM-L
Refer to caption Scene 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Meas GT Patch ADMM-Net CST-L+ DGSMP DIP-HSI HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DiffSCI $\lambda$-Net MST-L PnP-CASSI TSA-Net ASLE-SSM-L
Figure 4: Visual comparison on $\operatorname{KAIST}$. Top is $Scene$ 1 at wavelength 487.0nm. Bottom is $Scene$ 2 at wavelength 575.5nm
@@ 4.2 Simulation Results
Table 2: Comparisons of different frameworks on PSNR, SSIM, FLOPs, Params, and Inference Time on the test dataset.
Framework  MST-S mst  CST-S cst  ASLE-SSM-S
PSNR 34.14 34.53 34.74
SSIM 0.938 0.945 0.946
Params (M) 0.93 1.20 0.81
FLOPS (G) 12.96 11.67 7.01
Inference Time(s) 0.022 0.041 0.009
We conduct a comprehensive comparative analysis of our ASLE-SSM method and SOTA HSI reconstruction algorithms in Params, FLOPs, PSNR, and SSIM. We compared the results of ASLE-SSM with 12 SOTA methods, including two model-based methods (GAP-TV gap_tv , and DeSCI desci ), two CNN-based methods ($\lambda$-net lambda , TSA-Net tsa_net ), three Transformer-based methods (HDNet hdnet , MST-S, MST-M, MST-L mst , and CST-S, CST-M, CST-L, CST-L+ cst ), two Plug-and-Play (PnP) methods (DIP-HSI self , DiffSCI diffsci ) and three deep unfolding methods (ADMM-Net admm-net , GAP-Net gapnet , DGSMP gsm ) across 10 simulation scenes. The quantitative results of different models are listed in Table 1.
As shown in Table 1, our ASLE-SSM-L obtains 35.74dB of PSNR and 0.958 of SSIM, achieving the best reconstruction performance among SOTAs. When compared to CNN-based and Transformer-based methods such as TSA-net tsa_net , HDNet hdnet , MST-L mst , and CST-L+ cst , our ASLE-SSM-L demonstrates notable improvements over these methods, which are 4.39dB, 1.08dB, 0.67dB, and 0.03dB on average, respectively. The spectral density curves in Fig. 5 also demonstrate our superiority.
We also conduct comparisons of our ASLE-SSM-S with previous Transformer-based models MST mst and CST cst in the small size. As shown in Table 2, we not only achieve better performance in PSNR and SSIM compared to MST and CST but also obtain a more lightweight model with faster speed. Specifically, the ASLE-SSM-S is 2.4 times faster than MST-S and 4.5 times faster than CST-S, while having 0.12 (M) fewer parameters than the MST-S, which means the smallest parameter count. From Fig. 1, we can also observe that our ASLE-SSM not only achieves better results, but the value of FLOPs is also significantly lower than that of the corresponding MST and CST models.
Except for quantitative comparisons, we also visualize the results of simulation experiments. We choose one of 28 channels from Scene 1 and Scene 2, against the simulation results obtained from twelve SOTA approaches. Visually, as illustrated in Fig. 4, our method yields smoother and cleaner details. The experimental results indicate the capability of our method to generate high-quality HSIs. Specifically, in Scene 1, our reconstruction results better generate the curvature of the chin compared to other models. In Scene 2, our model clearly reconstructs the shapes of the blocks and ropes. The heart-like pattern on the blocks is also well-preserved.

### v2

@@ 3.1 Architecture of CS-Mamba
For the denoiser network, we employ a customized Mamba using a spatial-spectral SSM block as its fundamental unit. The architecture comprises an encoder, a bottleneck, and a decoder. Given a measurement $\mathbf{y}\in\mathbb{R}^{H(W+d(N_{\lambda}-1))}$, we reverse the dispersion and shift back $\mathbf{Y}$ to initialize the input signal $\mathbf{X}\in\mathbb{R}^{H\times W\times N_{\lambda}}$:
@@
We first concatenate the input $\mathbf{X_{0}}$ with the physical mask $\mathbf{Y}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ and then employ a 1$\times$1 convolution layer to integrate information from the original signal and the physical mask. Moreover, the integrated input $\mathbf{X_{0}}$ undergoes a 3$\times$3 convolution embedding into a feature $\mathbf{X_{0}^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times C}$. $\mathbf{X_{0}^{{}^{\prime}}}$ undergoes downsampling and upsampling before entering the decoder, followed by a 3$\times$3 convolution to extract residual information. Summing the integrated input and residual output yields the reconstructed HSIs. We adjust the number of spatial-spectral SSM blocks in the encoder, bottleneck, and decoder to create different CS-Mamba model sizes.
3.2 Enhanced Local SSM (LE-SSM)
Although Mamba effectively models long-range dependencies, directly using the SSM mechanism often limits its ability to capture local features. Recent methods integrate CNNs into the network to leverage their local feature extraction capabilities, enhancing original SSM performance. However, this increases the parameter count, undermining the lightweight nature of SSMs. While adding CNN can boost performance, it does not address the core issue. Inspired by LocalMamba [24], we propose a simplified local mechanism in the spectral dimension. This involves using two global scanning directions, traversing the image row by row in both forward and backward directions, while the other two scanning directions serve as local branches to unfold and scan local patches.
Figure 4: A visual comparison of with and w/o masked methods in our CS-Mamba model on real scene 1.
Table 1: Comparisons of our denoiser with previous End-to-End networks on PSNR, SSIM, FLOPs, Params, and Inference Time on the test dataset.
Framework  MST-S [3]  CST-S [2]  Cross-Scan SSM
PSNR 34.14 34.53 34.74
SSIM 0.938 0.945 0.946
Params (M) 0.93 1.20 0.81
FLOPS (G) 12.96 11.67 7.01
Inference Time(s) 0.022 0.041 0.009
The specific details of the scanning process are illustrated in Fig. 3(d). Given the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, we first generate four copies of input data with different arrangements. For the global branch, we directly flatten the feature map into $\mathbf{X_{1}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, and then reverse it to generate $\mathbf{X_{2}}$. For the local branch, we first set the size of our patch flattening each patch into a 1D sequence, and then linking each patch together to generate $\mathbf{X_{3}}$. Similarly, we reverse $\mathbf{X_{3}}$ to obtain $\mathbf{X_{4}}$. After the generation of input data in four directions, we conduct the process as follows:
$\mathbf{\hat{X}}=LE\text{-}SSM(\mathbf{X^{{}^{\prime}}}).$
The structure of LE-SSM is shown in Fig. 3(b). It is worth noting that our method does not employ any CNNs or spatial and channel attention mechanisms based on MLPs, ensuring that our method does not add any extra parameters or time consumption compared to the global cross-scan mechanism [28]. However, it also means we need to carefully choose the size of local patches, or else the reconstruction result will even become worse. In our implementation, we use patches of size 4$\times$4 to achieve a relatively better performance.
3.3 Cross-Scanning SSM (CS-SSM) Mechanism
Our LE-SSM mechanism creates a locally enhanced global receptive field but does not fully utilize the rich spectral information in HSIs. Recent methods [11, 43] have integrated channel information by scanning along the spectral dimension; however, this approach overlooks local features in both spectral and spatial dimensions. Neighboring pixels and bands often share similarities, but direct spectral scanning can increase the distance between highly correlated data points, reducing reconstruction quality.
Table 2: Comparisons of our CS-Mamba-S (9stg) with previous Transformer-based methods on PSNR, SSIM, FLOPs, and Params.
Framework  DAUHST [5]  DERNN [14]  CS-Mamba-S
PSNR 38.36 39.93 40.00
SSIM 0.967 0.976 0.980
Params (M) 6.15M 0.65M 0.57M
FLOPS (G) 79.50 81.99 65.77
To address the above issues, we propose a local scanning method that integrates local features from both spatial and spectral dimensions. This mechanism attempts to exploit the local properties of HSIs in both spatial and spectral dimensions while leveraging the rich spectral information to guide the reconstruction process. First, we rearrange the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ into patches along the spatial context by dividing the input HSIs into a series of smaller data cubes $\mathbf{S}\in\mathbb{R}^{P\times P\times C}$, which will be flattened into sequences for scanning. After that, we organize the scanning order within each sequence based on our spatial-spectral local cubes $\mathbf{C}\in\mathbb{R}^{h\times w\times c}$. As shown in Fig. 3(e), different colors indicate each spatial-spectral local cube. The scanning order of each data point is also depicted. The process can be formulated as follows:
$\mathbf{C}=Divide(\mathbf{X}),$ (16)
$\mathbf{C^{{}^{\prime}}}=CS\text{-}SSM(\mathbf{C}),$
where $\mathbf{C^{{}^{\prime}}}$ denotes the output patch of CS-SSM. After the SSM mechanism, we reorganize the patches back into the same order as the input $\mathbf{X}$. By scanning the data points within each data cube, we can leverage the local similarity between adjacent spectral bands and pixels, thus achieving better results than previous methods that directly scan along the spectral dimension.
@@ 3.3 Cross-Scanning SSM (CS-SSM) Mechanism
     Algorithms   Reference   Params   GFLOPs      S1      S2      S3      S4      S5      S6      S7      S8      S9      S10     Avg
DIP-HSI [34] ICCV 2021
@@
ADMM-Net [31] ICCV 2019
4.27M
78.58
34.03
0.919
33.57
0.904
34.82
0.933
39.46
0.971
31.83
0.924
32.47
0.926
32.01
0.898
30.49
0.907
33.38
0.917
30.55
0.899
33.26
0.920
DGSMP [23] TPAMI 2023
3.76M
646.65
33.26
0.915
32.09
0.898
33.06
0.925
40.54
0.964
28.86
0.882
33.08
0.937
30.74
0.886
31.55
0.923
31.66
0.911
31.44
0.925
32.63
0.917
DiffSCI [37] CVPR 2024
@@
DAUHST-9stg [5] NIPS 2022 6.15M 79.50
37.25
0.958
39.02
0.967
41.05
0.971
46.15
0.983
35.80
0.969
37.08
0.970
37.57
0.963
35.10
0.966
40.02
0.970
34.59
0.956
38.36
0.967
PADUT [26] ICCV 2023 5.38M 90.46
37.36
0.962
40.43
0.978
42.38
0.979
46.62
0.990
36.26
0.974
37.27
0.974
37.83
0.966
35.33
0.974
40.86
0.978
34.55
0.963
38.89
0.974
RDLUF [13] CVPR 2023 1.89M 115.34
37.94
0.966
40.95
0.977
43.25
0.979
47.83
0.990
37.11
@@
37.47
0.975
38.58
0.969
35.50
0.970
41.83
0.978
35.23
0.962
39.57
0.974
DERNN (9stg) [14] TGRS 2024 0.65M 81.99
38.26
0.965
40.97
0.979
43.22
0.979
48.10
0.991
38.08
0.980
37.41
0.975
38.83
0.971
36.41
0.973
42.87
0.981
35.15
0.962
39.93
0.976
DERNN (9stg)* [14] TGRS 2024 1.04M 134.18
38.49
0.968
41.27
0.980
43.97
0.980
48.61
0.992
38.29
0.981
37.81
0.977
39.30
0.973
36.51
0.974
43.38
0.983
35.61
0.966
40.33
0.977
SSR-L* [52] CVPR 2024 1.73M 78.93
38.81
0.968
41.51
0.979
43.76
0.979
48.62
0.988
38.32
0.979
37.85
0.975
38.50
0.969
36.85
0.974
42.64
0.980
35.82
0.965
40.27
0.976
CS-Mamba (3stg) Ours
0.97M
37.80
37.93
0.968
40.00
0.978
42.73
0.983
47.18
0.993
36.48
0.976
37.10
0.976
37.91
0.970
35.39
0.974
41.43
0.981
34.82
0.965
39.10
0.977
CS-Mamba (5stg) Ours
0.97M
62.93
38.23
0.971
40.62
0.982
43.44
0.984
47.58
0.994
37.95
0.982
37.74
0.979
38.71
0.974
36.27
0.979
42.89
0.986
35.45
0.970
39.90
0.980
CS-Mamba (9stg) Ours
0.97M
113.18
38.51
0.973
41.30
0.984
43.77
0.985
47.55
0.994
38.72
0.985
38.05
0.981
38.88
0.976
36.64
0.981
43.06
0.987
@@
43.25
0.988
35.78
0.971
40.42
0.983
CS-Mamba-s (9stg) Ours
0.57M
65.77
38.63
0.972
40.90
0.983
43.46
0.984
47.30
0.994
38.41
0.983
37.68
0.979
38.80
0.975
36.29
0.979
43.09
0.986
35.41
0.969
40.00
0.980
Figure 5: A visual comparison of the performance of MST, CST, and DAUHST methods in the real scene.
3.4 Masked Training Strategy
Before presenting our masked training strategy, we discuss our observed phenomena and motivation for this approach. In Fig. 5, we display visual results for three Transformer-based methods [3, 2, 5] in a real scene. While their simulation performance improves from left to right, visual quality degrades in real scenarios. Compared to MST, CST’s reconstructed image shows more noise artifacts on the nose and forehead. These spots, absent in ground truth, suggest noise from 2D compressed measurements. DAUHST’s results exhibit even greater degradation, which is consistently observed across other algorithms in this scene.
This effect stems from two factors. First, simulation-based image reconstruction, aimed at minimizing compression loss, struggles with unknown real-world noise distributions. While shot noise is added during training, it inadequately reflects real-world complexities, leading to insufficient generalization. Secondly, focusing on mask pattern recovery can overemphasize specific information, weakening resilience to other noise types, and creating a balance challenge between compression and real-world losses.
$Total\ Loss=Compression\ Loss+\epsilon,$ (17)
where $\epsilon$ represents the complex real noise.
We propose a masked training strategy inspired by [8] for SCI reconstruction to enhance robustness and generalization. During training, a randomly generated 0-1 mask is applied to feature maps, creating an inpainting task that compels the network to reconstruct occluded regions. This strengthens the model’s capacity to handle complex real-world scenarios. Unlike traditional image denoising, where noise is added to clean images, SCI reconstruction starts with compressed 2D measurements that inherently carry compression loss, compounded by mixed noise in practice. By consistently applying the same mask during both training and testing, we ensure alignment between phases, leading to improved reconstruction accuracy and visually appealing results. This approach addresses unique SCI challenges, enhancing model performance under varied real-world conditions.
@@ 4.1 Experimental Settings
The effectiveness of our CS-Mamba network is verified on both simulation and real datasets, by using 28 wavelengths ranging from 450nm to 650nm, acquired via spectral interpolation manipulation.
Simulation HSI Data. Two simulated HSI datasets CAVE [38] and KAIST [10] are utilized to conduct the simulation experiment. The CAVE dataset contains 32 HSIs with a spatial size of $512\times 512$, while the KAIST dataset includes 30 HSIs with a spatial size of $2704\times 3376$. Following the experimental settings of TSA-Net [32], we designate CAVE as the training set and select 10 scenes from the KAIST as the testing set.
Real HSI Data. We adopt the real HSIs obtained by the CASSI system, as developed in TSA-Net [32] for testing.
Evaluation Metrics. We select peak signal-to-noise ratio (PSNR) [7] and structural similarity (SSIM) [46] as the evaluation metrics to assess the quality of HSI reconstruction.
Implementation Details. Our CS-Mamba is implemented using PyTorch and we conduct model training with Adam optimizer ($\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999) for 300 epochs. The learning rate is set to $1\times 10^{-3}$ at the beginning and a cosine annealing scheduler is adopted. Image patches with size $256\times 256$ are randomly cropped from the 3D HSI cubes with 28 channels for training in the simulation and real experiments. The shifting step $d$ in the dispersion process is set to 2. We apply the data augmentation techniques including random flipping and rotation. Models are trained on RTX 4090 GPUs.
Refer to caption Scene 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Meas GT Patch ADMM-Net CST-L+ DGSMP DIP-HSI HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DiffSCI $\lambda$-Net MST-L DAUHST PADUT CS-Mamba
Figure 6: Visual comparison of different methods on simulation $Scene$ 2 at wavelength 575.5nm
@@ 4.2 Simulation Results
We conduct a comprehensive comparative analysis of our CS-Mamba method and SOTA HSI reconstruction algorithms in Params, FLOPs, PSNR, and SSIM. We compared the results of CS-Mamba with 11 SOTA methods algorithms across 10 simulation scenes. The quantitative results of different models are listed in Table 3.
As shown in Table 3, our CS-Mamba-9stg* obtains 40.42dB of PSNR and 0.983 of SSIM, achieving the best reconstruction performance among SOTAs. When compared to recent Transformer-based methods such as DAUHST [5], RDLUF [13], DERNN [14], and SSR-L* [52], our CS-Mamba demonstrates notable improvements over these methods, which are 2.06dB, 0.85dB, 0.49dB, and 0.15dB on average, respectively.
We also conduct comparisons of our CS-Mamba-S with previous Transformer-based methods DAUHST [5] and DERNN [14]. As shown in Table 2, we not only achieve better performance in PSNR and SSIM compared to DAUHST and DERNN but also obtain a more lightweight model. Specifically, the CS-Mamba has 0.08 (M) fewer parameters than the DERNN and 13.73 (G) fewer in FLOPs, which means the smallest parameter count and Computational cost among recent deep unfolding methods. Furthermore, we evaluated the denoiser network (Cross-Scan SSM Block) as an independent end-to-end model, comparing it against previous end-to-end Transformer methods, as presented in Table 1. Specifically, the Spatial-Spectral SSM is 2.4 times faster than MST-S and 4.5 times faster than CST-S, while having 0.12 (M) fewer parameters than the MST-S, which means the smallest parameter count.
Except for quantitative comparisons, we also visualize the results of simulation experiments. We choose one of 28 channels from Scene 2, against the simulation results obtained from eleven SOTA approaches. Visually, as illustrated in Fig. 6, our method yields smoother and cleaner details. The experimental results indicate the capability of our method to generate high-quality HSIs.